OpenStudy (anonymous):

The regular price of an item is $84. The item is on sale with a discount of $12. What is the discount rate?

OpenStudy (mathstudent55):

discount rate = (\ \dfrac{discount}{original ~price} \times 100 \)

OpenStudy (anonymous):

what? lol

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Divide the discount by the original price and multiply by 100. That will give you the discount rate as a percent.

OpenStudy (anonymous):

14.29

OpenStudy (mathstudent55):

14.29%

OpenStudy (mathstudent55):

\(discount ~rate = \dfrac{discount}{original ~price} \times 100 \)